If a 1.00 kg body has an acceleration of 9.64 m/s2 at 58 to the positive direction of the x axis, then what are (a) the x component and (b) the y component of the net force on it, and (c) what is the net force in unit-vector notation? (a) Number Units (b) Number Units j Units (c) Number
Answer #1

(a)

From the Newton second law of motion , the net force acting on the body is

F = ma = 1.0 kg ( 9.64 m/s^2) = 9.64 N

x component of force

Fx = F cos theta = 9.64 N cos 58 = 5.10 N

(b)

y component is

Fy = F sin theta = 9.64 N sin 58 = 8.17 N

(c)

F = 5.10 N i + 8.17 N j

magnitude

F = sqrt ( 5.10)^2 + (8.17)^2 = 9.64 N